# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ 3. Conditioned limit theorems for Lévy processes
Take $\gamma\in(0,\frac{1}{2})$ and fix an $N>0$. For $t>N$, we have
(4.41) $Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(x,t)$ $\leq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\}}e^{-\int_{t-N}^{t}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\xi_{s},t-s))\mathrm{d}s}\right]$
(4.42) $\leq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right]$
(4.43) $=:e^{-\alpha t}(J_{1}(t)+J_{2}(t)),$
@@
$J_{1}(t):=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}\geq\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right],$
$J_{2}(t):=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right].$
Since $Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(x,t)$ is increasing in $x$, we have
(4.44) $J_{1}(t)\leq e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma},s))\mathrm{d}s}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\right).$
By (4.15) and (1.12),
(4.45) $Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(x,t)\geq g(t)\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\right).$
Thus,
(4.46) $e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma},s))\mathrm{d}s}$
(4.47) $\leq$ $\exp\Big{\{}-\int_{0}^{N}\varphi\left(g(s)\mathbf{P}_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}}\left(\tau_{0}^{-}>s,\xi_{s}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\right)\right)\Big{\}}.$
[…]
@@
(4.49) $\limsup_{N\to\infty}\limsup_{t\to\infty}J_{1}(t)\leq C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x)\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
Next, we show that $\lim_{t\to\infty}J_{2}(t)=0$. For $\epsilon>0$, it holds that
$J_{2}(t)\leq$ $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}\right)$
$\leq$ $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t<\xi_{t}\leq\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t\right)$
$+\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}\right).$
@@
(4.50) $\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t<\xi_{t}\leq\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t\right)$
(4.51) $=\frac{\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x)}{\sqrt{2\pi}}\int_{\frac{y}{\sigma}}^{\frac{y+\epsilon}{\sigma}}e^{-\frac{z^{2}}{2}}\mathrm{d}z\xrightarrow{\epsilon\to 0}0.$
[…]
(4.52) $\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}\right)$
(4.53) $\leq\mathbf{P}_{x}\left(A_{t}\right)+\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma},A_{t}^{c}\right)$
By Lemma 3.2, we have $\lim_{t\to\infty}\mathbf{P}_{x}\left(A_{t}\right)=0$. Furthermore, using the reflection principle for Brownian motion, we get that, as $t\to\infty$,
$\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma},A_{t}^{c}\right)$
$\leq$ $\mathbf{Q}_{x}\left(W_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left[\xi_{1}\right]t-t^{\frac{1}{2}-\kappa},\inf_{r\in[t-N,t]}W_{r}<\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t+t^{\gamma}+t^{\frac{1}{2}-\kappa}\right)$
@@
Fix an $N>0$. Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1) and (H2). If $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, then we have
(4.54) $\lim_{t\to\infty}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right]$
@@
(4.56) $\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right]$
(4.57) $=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t-N\}}\mathbf{E}_{\xi_{t-N}}\left[1_{\{\tau_{0}^{-}>N,\xi_{N}>y\}}e^{-\int_{0}^{N}\varphi(Q_{y}(\xi_{s},N-s))\mathrm{d}s}\right]\right]$
(4.58) $=:\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t-N\}}f_{N}^{y}(\xi_{t-N})\right],$
@@
(4.59) $f_{N}^{y}(z):=\mathbf{E}_{z}\left[1_{\{\tau_{0}^{-}>N,\xi_{N}>y\}}e^{-\int_{0}^{N}\varphi(Q_{y}(\xi_{s},N-s))\mathrm{d}s}\right].$
@@
which implies that $f_{N}^{y}(z)$ is bounded and increasing with respect to $z$. Then $f_{N}^{y}$ is a.e. continuous. Applying Theorem 3.5 to $f_{N}^{y}$, we get that
(4.61) $\lim_{t\to\infty}t^{3/2}e^{-\Psi(\lambda_{*})t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right]$
@@
(4.63) $Q_{y}(x,t)\leq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right].$
@@
Moreover, using the fact that $Q_{y}(x,t)\leq g(t)$ and Lemma 4.2, we get
(4.66) $Q_{y}(x,t)\geq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right]e^{-\int_{0}^{t-N}\varphi(g(t-s))\mathrm{d}s}.$
@@
(4.74) $\lim_{t\to\infty}\mathbb{P}_{x}\left(M_{t}\in(a,b]|\zeta>t\right)=\lim_{t\to\infty}\frac{Q_{a}(x,t)-Q_{b}(x,t)}{u(x,t)}=\frac{\int_{a}^{b}e^{-\lambda_{*}z}\widehat{R}^{*}(z)\mathrm{d}z}{\int_{0}^{\infty}e^{-\lambda_{*}z}\widehat{R}^{*}(z)\mathrm{d}z}.$
@@
$\mathbb{P}_{x}\left(M_{t}>y|\zeta>t\right)=\frac{Q_{y}(x,t)}{u(x,t)}$
$\leq\frac{e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right)}{C_{sub}e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t\right)}=\frac{1}{C_{sub}}\mathbf{P}_{x}\left(\xi_{t}>y|\tau_{0}^{-}>t\right).$
@@
For any $0<x<y$, define
@@
(5.2) $v(x,y)=\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\alpha\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right],$
@@
(5.3) $v(x,y)=\mathbf{E}_{x}\left[1_{\{\tau_{z}^{+}<\tau_{0}^{-}\}}v(\xi_{\tau_{z}^{+}},y)e^{-\alpha\tau_{z}^{+}-\int_{0}^{\tau_{z}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right].$
@@
$+\int_{0}^{\infty}\beta e^{-\beta s}\mathbf{E}_{x}\left(\left(1-\sum_{k=0}^{\infty}p_{k}\left(1-v(\xi_{s},y)\right)^{k}\right)1_{\{\tau_{y}^{+}\land\tau_{0}^{-}>s\}}\right)$
@@
By [4, Lemma 4.1], the above equation is equivalent to
@@
where $\Phi$ is defined in (1.6). Using the Feynman-Kac formula, we get that
(5.5) $v(x,y)=\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\alpha\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$
This gives the desired result. ∎
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process. For any $0<x<y$, we have
@@
(5.7) $v(x,y)\leq\mathbf{E}_{x}\left[e^{-\alpha\tau_{y}^{+}}1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}\right]=\frac{W^{(\alpha)}(x)}{W^{(\alpha)}(y)},\quad x<y.$
@@
Proof of Theorem 1.4: By Lemmas 5.1 and 2.4, we have
$v(x,y)$ $=\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\alpha\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]$
$=e^{(x-y)\psi(\alpha)}\mathbf{E}_{x}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right].$
@@
(5.9) $v(x,y)$ $:=e^{(x-y)\psi(\alpha)}\mathbf{E}_{x}^{\psi(\alpha)}\left[1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]$
(5.10) ×$\qquad\times\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]=:e^{(x-y)\psi(\alpha)}A_{1}(x,y)A_{2}(y),$
@@
$A_{1}(x,y):=\mathbf{E}_{x}^{\psi(\alpha)}\left[1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right],$
$A_{2}(y):=\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right].$

### v2

@@ 3. Conditioned limit theorems for Lévy processes
Take $\gamma\in(0,\frac{1}{2})$ and fix an $N>0$. For $t>N$, using Lemma 4.2, we have
(4.52) $Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(x,t)$ $\leq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t\}}e^{-\int_{t-N}^{t}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(\xi_{s},t-s))\mathrm{d}s}\right)$
(4.53) $\leq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right)$
(4.54) $=:e^{-\alpha t}(K_{1}(t)+K_{2}(t)),$
@@
$K_{1}(t):=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}\geq\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right),$
$K_{2}(t):=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right).$
Repeating the argument leading to (4.36), we obtain that
(4.55) $\limsup_{N\to\infty}\limsup_{t\to\infty}\frac{K_{1}(t)}{\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t\right)}\leq C_{sub}.$
@@
(4.56) $\limsup_{N\to\infty}\limsup_{t\to\infty}K_{1}(t)\leq q_{x}C_{sub}\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
Next, we show that $\lim_{t\to\infty}K_{2}(t)=0$. For $\epsilon>0$, it holds that
$K_{2}(t)\leq$ $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma}\right)$
$\leq$ $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t<\xi_{t}\leq\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t\right)$
$+\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma}\right).$
@@
(4.57) $\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t<\xi_{t}\leq\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t\right)\xrightarrow{\epsilon\to 0}0.$
For $\mathbf{E}_{0}(\xi_{1})>0$, since $\left((\xi_{t}-\mathbf{E}_{0}(\xi_{1})t)_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}}\right)$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1}-\mathbf{E}_{0}(\xi_{1}))=0$, it follows from Lemma 3.2 that there exists a Brownian motion $W$ with diffusion coefficient $\sigma^{2}=\mathbf{E}_{0}(\xi_{1}^{2})$ starting from the origin such that for all $t\geq 1$,
(4.58) $\mathbf{P}_{x}\left(\sup_{s\in[0,1]}|(\xi_{ts}-\mathbf{E}_{0}(\xi_{1})ts)-x-W_{ts}|>t^{\frac{1}{2}-\kappa}\right)\leq\frac{C_{3}(\kappa)}{t^{(\frac{1}{2}-\kappa)(\delta+2)-1}},$
where $\kappa$ and $C_{3}(\kappa)$ are defined in Lemma 3.2. Let
(4.59) $D_{t}:=\left\{\sup_{s\in[0,1]}|(\xi_{ts}-\mathbf{E}_{0}(\xi_{1})ts)-x-W_{ts}|>t^{\frac{1}{2}-\kappa}\right\},$
then by (4.58), we have $\lim_{t\to\infty}\mathbf{P}_{x}\left(D_{t}\right)=0$. Moreover, we have
(4.60) $\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma}\right)$
(4.61) $\leq\mathbf{P}_{x}\left(D_{t}\right)+\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma},D_{t}^{c}\right).$
Furthermore, using the reflection principle for Brownian motion, we get that, as $t\to\infty$,
$\mathbf{P}_{x}\left(\xi_{t}>\sqrt{t}(y+\epsilon)+\mathbf{E}_{0}\left(\xi_{1}\right)t,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t+t^{\gamma},D_{t}^{c}\right)$
$\leq$ $\mathbf{Q}_{x}\left(W_{t}>\sqrt{t}(y+\epsilon)-t^{\frac{1}{2}-\kappa},\inf_{r\in[t-N,t]}W_{r}<\sqrt{t}y+t^{\gamma}+t^{\frac{1}{2}-\kappa}\right)$
@@
Fix an $N>0$. Assume that (1.7) holds and $\xi$ is a Lévy process satisfying (H1), (H2) and (H3). If $\mathbf{E}_{0}\left(\xi_{1}\right)<0$, then we have
(4.62) $\lim_{t\to\infty}t^{3/2}e^{-\Psi(\lambda_{*})t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right)$
@@
(4.64) $\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right)$
(4.65) $=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t-N\}}\mathbf{E}_{\xi_{t-N}}\left(1_{\{\tau_{0}^{-}>N,\xi_{N}>y\}}e^{-\int_{0}^{N}\varphi(Q_{y}(\xi_{s},N-s))\mathrm{d}s}\right)\right)$
(4.66) $=:\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t-N\}}f_{N}^{y}(\xi_{t-N})\right),$
@@
(4.67) $f_{N}^{y}(z):=\mathbf{E}_{z}\left(1_{\{\tau_{0}^{-}>N,\xi_{N}>y\}}e^{-\int_{0}^{N}\varphi(Q_{y}(\xi_{s},N-s))\mathrm{d}s}\right).$
@@
which implies that $f_{N}^{y}(z)$ is bounded and increasing with respect to $z$. Then $f_{N}^{y}$ is a.e. continuous. By [9, Corollary 3.2], $f_{N}^{y}(z)e^{-\lambda_{*}z}(1+|z|)$ is directly Riemann integrable with respect to $z$. Applying Theorem 3.5 with $f$ replaced by $f_{N}^{y}$, we get
(4.69) $\lim_{t\to\infty}t^{3/2}e^{-\Psi(\lambda_{*})t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right)$
@@
(4.71) $Q_{y}(x,t)\leq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right).$
@@
Moreover, using the fact that $Q_{y}(x,t)\leq g(t)=\mathbb{P}_{0}(\widetilde{\zeta}>t)$ and Lemma 4.2, we get
(4.74) $Q_{y}(x,t)\geq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{t-N}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right)e^{-\int_{0}^{t-N}\varphi(g(t-s))\mathrm{d}s}.$
@@
(4.82) $\lim_{t\to\infty}\mathbb{P}_{x}\left(M_{t}\in(a,b]|\zeta>t\right)=\lim_{t\to\infty}\frac{Q_{a}(x,t)-Q_{b}(x,t)}{u(x,t)}$
(4.83) $=$ $\frac{\lim_{N\to\infty}\int_{0}^{\infty}\mathbb{P}_{z}(M_{N}\in(a,b])e^{-\lambda_{*}z}\widehat{R}^{*}(z)\mathrm{d}z}{\lim_{N\to\infty}\int_{0}^{\infty}\mathbb{P}_{z}(M_{N}\in(0,\infty))e^{-\lambda_{*}z}\widehat{R}^{*}(z)\mathrm{d}z}.$
@@
$\mathbb{P}_{x}\left(M_{t}>y|\zeta>t\right)=\frac{Q_{y}(x,t)}{u(x,t)}\leq\frac{e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right)}{C_{sub}e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t\right)}=\frac{1}{C_{sub}}\mathbf{P}_{x}\left(\xi_{t}>y|\tau_{0}^{-}>t\right).$
@@
Recall that $\alpha=\beta(1-m)$. For any $0<x<y$, define
@@
(5.2) $v(x,y)=\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\alpha\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right),$
@@
(5.3) $v(x,y)=\mathbf{E}_{x}\left(1_{\{\tau_{z}^{+}<\tau_{0}^{-}\}}v(\xi_{\tau_{z}^{+}},y)e^{-\alpha\tau_{z}^{+}-\int_{0}^{\tau_{z}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$
@@
$+\int_{0}^{\infty}\beta e^{-\beta s}\mathbf{E}_{x}\left(\left(1-\sum_{k=0}^{\infty}p_{k}\left(1-v(\xi_{s},y)\right)^{k}\right)1_{\{\tau_{y}^{+}\land\tau_{0}^{-}>s\}}\right)\mathrm{d}s$
@@
By [5, Lemma 4.1], the above equation is equivalent to
@@
where $\Phi$ is defined in (1.6). By repeating the argument leading to (4.7), we get the desired result. ∎
In the remainder of this section, we always assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process.
@@
Assume that $\xi$ is a spectrally negative Lévy process. For any $0<x<y$, we have
@@
(5.6) $v(x,y)\leq\mathbf{E}_{x}\left(e^{-\alpha\tau_{y}^{+}}1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}\right)=\frac{W^{(\alpha)}(x)}{W^{(\alpha)}(y)},\quad x<y.$
@@
Proof of Theorem 1.4: By Lemmas 5.1 and 2.3, we have
$v(x,y)$ $=\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\alpha\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)$
$=e^{(x-y)\psi(\alpha)}\mathbf{E}_{x}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$
@@
(5.8) $v(x,y)$ $=e^{(x-y)\psi(\alpha)}\mathbf{E}_{x}^{\psi(\alpha)}\left(1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)$
(5.9) $\qquad\times\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)=:e^{(x-y)\psi(\alpha)}A_{1}(x,y)A_{2}(y),$
@@
$A_{1}(x,y):=\mathbf{E}_{x}^{\psi(\alpha)}\left(1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right),$
$A_{2}(y):=\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$